OPTICAL ELECTRON PARAMETER SEMICONDUCTING NANOSTRUCTURES

N. P. NETESOVA, M. V. Lomonosov Moscow State University, Physics Faculty, 119992, Moscow, Sparrow Hills, Russia.

Problem: to not use the classical Kramers-Kronig integral transformation and to define all optical electron oscillation parameters for any energy point from semiconducting nanostructure experimental reflection spectra^{a}. Within the untied oscillation model the calculation technique of all semiconducting heterostructure optical parameters by the intermediate functions (h̄ · ω_{π}, h̄ · ω_{n}, h̄ · Γ) are the plasma, effective natural, radiant friction energies in eV, 2 · π · h̄ is the Planck constant) is presented. As an example the optical parameters of PbS, PbSe, PbTe and GaAs, GaP between 0 and 25 eV in any spectrum region are established. The consistent approximation approach of the reflectance factor R to real value is advanced. As a result, all heterostructure basic electron optical functions (h̄ · ω_{p}, h̄ · ω_{pm}, h̄ · ω_{c}, h̄ · γ are the plasma, plasma maximum, effective natural, radiant friction energies, ε_{r}, ε_{ι}, n_{r}, n_{ι} are the real and imaginary components of the dielectric ε and refractive index n functions, accordingly, (ε_{r})_{max}, (ε_{r})_{min}, (h̄ · ω) • ε_{ι} is conductivity, (h̄ · ω) • n_{ι} = (c · h̄/2) • α, where c is the light velocity, α is absorption coefficient, L = Im(−1/ε) are electron lossis, equal imaginary component of the minus reciprocal dielectric function ε, h̄ · ω • L = (h̄ · ω) • Im(−1/ε) are effective electron lossis) calculated by the intermediate functions in any electron optical spectrum region. Then, for GaP experimental reflection spectra it is selected the point h̄^{2} ·ω^{2} = 10.5625•10^{−4}, the intermediate parameters are h̄^{2} ·ω_{π}^{2} = 10.5625•10^{−4}, h̄^{2} · ω_{n}^{2} = 9.03130933157 • 10^{−4}, h̄^{2} · Γ^{2} = 1.875029665786 • 10^{−4}, the basic parameters are h̄^{2} · ω_{p}^{2} = 19.5902684716 • 10^{−4}, h̄^{2} · ω_{c}^{2} = 5.28479085993 • 10^{−4}, h̄^{2} · γ^{2} = 0.79237637701 • 10^{−4} (eV)^{2}. The R values calculated by electron parameters coincide with the experimental values R(h̄ω) to within 10^{−6} ÷10^{−10} for 12 symbol computation. By presented method the nanostructure oscillation electron parameters are determined for device producing.

aN. P. Netesova, NGS12 Proceedings, Editors: J. Kono, Jean Leotin, Toulouse, France,2, 178-183, (3-7 July 2005)